OpenStudy (anonymous):

Find the quotient of 2^3 divided by 5^2

OpenStudy (kamibug):

\(\huge\cal\color{indigo}{Welcome~to~Openstudy!~:)}\) 2^3 = 2*2*2 = ? 5^2 = 5*5 = ?